# Density-valued time series: Nonparametric density-on-density regression

### v1

@@ 2.2.3 Dependence model
To better understand the dependence structure of our data, our framework combines two random mechanisms: 1) the underlying process {}$\{\mathscr{F}\}_{t\in\mathbb{Z}}$ generating the random PDFs $\{f_{1},\ldots,f_{N+1}\}$ and 2) the random process providing the $N+1$ samples $\left\{X_{t,1},\ldots,X_{t,n_{t}}\right\}_{t=1,\ldots,N+1}$ for estimating the non-observable random PDFs. For any $t=1,2,\ldots,N$, set $\bm{X}_{t}=\left\{X_{t,1},\ldots,X_{t,n_{t}}\right\}$; the scheme hereafter sums up the overall random situation:
Density-valued process $f_{1}$longrightarrow $f_{2}=m(f_{1})\oplus\epsilon_{1}$longrightarrow $\cdots$ $\longrightarrow$ $f_{N+1}=m(f_{N})\oplus\epsilon_{N}$
[…]
@@
[…]
Estimated densities $\widehat{f}_{1}$ $\widehat{f}_{2}=\widehat{m}_{N}(\widehat{f}_{1})$ $\widehat{f}_{N+1}=\widehat{m}_{N}(\widehat{f}_{N})$
We assume that the random process of two consecutive densities $\{(\mathscr{F}_{t},\mathscr{F}_{t+1})\}_{t\in\mathbb{Z}}$ is a strictly stationary $\rho$-mixing process (see, e.g., Rio, 1993; Bradley, 2005). We first remind the definition of the $\rho$-mixing dependence condition. For any $\sigma$-field $\mathcal{A}$, let $\mathcal{L}^{2}(\mathcal{A})$ denote the space of square-integrable $\mathcal{A}$-measurable real-valued random variables. Given two $\sigma$-fields $\mathcal{A}$ and $\mathcal{B}$, $\rho(\mathcal{A},\mathcal{B})=\sup\big{\{}\left|\text{Corr}(U,\,V)\right|$, $\,U\in\mathcal{L}^{2}(\mathcal{A}),\,V\in\mathcal{L}^{2}(\mathcal{B})\big{\}}$ measures the dependence between $\mathcal{A}$ and $\mathcal{B}$ in terms of correlation. $\rho$ is between 0 and 1; the closer this quantity is to 0, the weaker the dependency. Let $\{\mathscr{Z}_{t}\}_{t\in\mathbb{Z}}$ be a random process and for any $k<\ell$, define $\mathcal{A}_{k}^{\ell}$ as the $\sigma$-filed generated by the $\mathscr{Z}_{t}$’s for $t$ between $k$ and $\ell$. The random sequence $\{\mathscr{Z}_{t}\}_{t\in\mathbb{Z}}$ is said to be $\rho$-mixing if the dependence coefficients $\rho(m)=\sup_{j\in\mathbb{Z}}\rho\left(\mathcal{A}_{-\infty}^{j},\,\mathcal{A}_{j+m}^{+\infty}\right)$ tends to 0 as $m$ goes to infinity. This asymptotic property is usually called the $\rho$-mixing condition. Assuming that $\{(\mathscr{F}_{t},\mathscr{F}_{t+1})\}_{t\in\mathbb{Z}}$ is $\rho$-mixing is equivalent to suppose that its dependence coefficients fulfill the $\rho$-mixing condition with $\mathscr{Z}_{t}=(\mathscr{F}_{t},\mathscr{F}_{t+1})$. Then, according to the scheme generating the data, it is reasonable to suppose that the random sequences $(f_{t},\,f_{t+1})_{t=1,2,\dots}$ and $(\bm{X}_{t},\,\bm{X}_{t+1})_{t=1,2,\dots}$ inherit the structure (i.e strictly stationary and $\rho$-mixing) of $\{(\mathscr{F}_{t},\mathscr{F}_{t+1})\}_{t\in\mathbb{Z}}$. As a by-product, the random sequence $(\widehat{f}_{t},\,\widehat{f}_{t+1})_{t=1,2,\dots}$ based on the estimated densities is also a strictly stationary $\rho$-mixing process. Lastly, to complete the structure of our data, conditionally to $f_{t}$ and for each $t=1,\ldots,N$, the error model $\epsilon_{t}$ is assumed independent of $\bm{X}_{t}$.
@@ 2.3 Asymptotic properties
This section is devoted to the theoretical properties of our density’s functional kernel estimator. We first introduce notations and assumptions we need to derive the asymptotic behaviour of $\widehat{m}(f)$.
@@
    It exists $C>0$ such that, for all $f$ and $g$ in $\mathcal{B}^{2}(I)$, $\|m(f)\ominus m(g)\|_{\mathcal{B}}\,\leq\,C\,\|f\ominus g\|_{\mathcal{B}}$,
@@
    it exists $a>1$ such that the $\rho$-mixing coefficient fulfills the condition $\rho(k)=O(k^{-a})$,
@@
    set $n=\inf_{t\geq 1}n_{t}$ and let $n=n_{N}$ be a sequence depending on $N$ such that $n$ goes to infinity with $N$; one assumes the existence of a sequence $\delta_{N}$ tending to 0 as $N$ goes to infinity such that, for any $t$, $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|^{2}_{\mathcal{B}}\right)=O(\delta_{N}^{2})$,
@@
    For any deterministic $f$ in $\mathcal{B}^{2}(I)$, set $\pi_{f}(h_{r})=P\left(\|f_{1}\ominus f\|_{\mathcal{B}}<h_{r}\right)$; it exists $C>0$ and $C^{\prime}>0$, $0<C\pi_{f}(h_{r})\leq\pi_{f}\left(h_{r}+o(h_{r})\right)\leq C^{\prime}\pi_{f}(h_{r})$,
@@
    $h_{r}$ tends to 0 with $N$, $N^{a/(1+a)}\,\pi_{f}(h_{r})$ goes to infinity with $N$, and it exists $b$ with $11/3<b<$ 1 such that $\delta_{N}^{2b}=o\left(\pi_{f}(h_{r})\right)$ and $\delta_{N}^{1-b}=o\left(h_{r}\right)$,
@@
[…]
(H(H1)) is usual; it requires that the regression operator is smooth enough by fulfilling the Lipschitz property. In (H(H3)), the model dependency assumes an arithmetic decay for the mixing coefficients. This assumption simplifies the convergence rate, but more general writing involving the mixing coefficients $\rho(N)$ can be obtained. (H(H5)) relates, for each $t$, the size $n_{t}$ of the sample $X_{t,1},\ldots,X_{t,n_{t}}$ to the number $N$ of random PDFs. In addition, it requires that the mean integrated squared error $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)$ of $\widehat{f}_{t}$ tends to zero when $n_{t}$ (i.e. $N$) goes to infinity. In the standard multivariate setting, the sequence $\delta_{N}$ can be interpreted as an upper bound of the well-known quantity $\text{MISE}(\widehat{f}_{t})=\text{E}\left(\|\widehat{f}_{t}-f_{t}\|\right)$ (for more details on MISE, see Rosenblatt, 1956; Epanechnikov, 1969; Wegman, 1972; Nadaraya, 1974; Marron & Wand, 1992; Hansen, 2005, among others). It is shown Section A in Appendix that as soon as $f\in\mathcal{B}^{2}(I)$ with $f\geq\theta>0$, $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)=O\left\{\text{MISE}(\widehat{f}_{t})\right\}$. In other words, when the density $f$ admits a strictly nonnegative lower bound, then $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)$ inherits the asymptotic properties of MISE. The hypothesis (H(H7)) assumes that the behaviour of the small ball probability $\pi_{f}(h)$ is not sensitive to variations on radius negligible with respect to the bandwidth $h$ (see Li & Shao, 2001 for a survey on small ball probability). According to the isometry property of the clr transformation (see Section 2.1), $\pi_{f}(h_{r})=P(\|X-x\|_{2}<h_{r})$ with $X=\text{clr}(f_{1})$ and $x=\text{clr}(f)$ in $\mathcal{L}^{2}_{0}(I)$; the class of $\mathcal{L}^{2}_{0}(I)$-valued random functions fulfilling (H(H7)) encompasses Gaussian-type random processes (see Ferraty et al., 2019). (H(H9)) relates the asymptotic behaviour of $h_{r}$, $\pi_{f}(h_{r})$, $N$ and $\delta_{N}^{2}$, the upper bound of $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)$ for any $t$. As $\pi_{f}(h_{r})$ tends to zero with $N$ faster than $h_{r}$ does, it is reasonable to require that $02b>1-b>$ 0 which results in $11/3<b<$ 1. With (H(H11)) and for any $q>0$, $\text{E}\left\{K^{q}_{r}(h^{-1}\,\|f_{t}\ominus f\|_{\mathcal{B}})\right\}$ is bounded below and above by a quantity of order $\pi_{f}(h_{r})$. For simplicity, we use a basic hypothesis on the kernel function. Nevertheless, it is possible to get the more accurate result $\text{E}\left\{K^{q}_{r}(h_{r}^{-1}\,\|f_{t}\ominus f\|_{\mathcal{B}})\right\}\,=\,C_{f,q}\pi_{f}(h_{r})\{1+o(h_{r})\}$ as soon as the kernel function $K_{r}$ is continuously differentiable on $(0,1)$ with $K_{r}(1)>0$, and for all $s\in(0,1)$, $K_{r}^{\prime}(s)\leq 0$ and the ratio $\pi_{f}(h_{r}\,s)/\pi_{f}(h_{r})$ tends to some positive value as $h$ tends to 0.
Once notations and assumptions are introduced, one can provide our main result by giving the pointwise convergence rate of our density-on-density kernel regression estimator; the proof is postponed to Section B of the Appendix.
@@
If (H(H1))-(H(H11)) are fulfilled, then
$\left\|\widehat{m}_{N}(f)\ominus m(f)\right\|_{\mathcal{B}}\,=\,O(h_{r})+O\left(\delta_{N}\,\pi_{f}(h_{r})^{-1/2}\right)+O_{P}\left(\left\{N^{a/(1+a)}\,\pi_{f}(h_{r})\right\}^{-1/2}\right).$
The first two terms correspond to the bias part. $O(h_{r})$, which is not surprising (see Ferraty & Vieu, 2006), is derived from the regularity assumption acting on the regression operator $m$ (see (H(H1))). The second term, which is not usual, involves $\delta_{N}$ and the small ball probability $\pi_{f}(h_{r})$. This is the price to pay when the random PDFs are not directly observed and are approximated by means of some estimating procedure, where $\delta_{N}$ can be interpreted as the upper bound of the root MISE of $\widehat{f}_{t}$ for any $t$. For the variance part, remember that $a$ controls the asymptotic behaviour of the mixing coefficient. So, the impact of the dependency model appears directly in the exponent of the sample size $N$ of the random densities. In the situation of a sample of independent explanatory functional random variables, we would get $N$ instead of $N^{a/(1+a)}$ (see again Ferraty & Vieu, 2006). As expected, the dependence structure of the data degrades the rate of convergence of $\widehat{m}$.
Let us now focus on the estimation of the random density $f_{t}$ with a particular attention on the sequence $\delta_{N}$ which is an upper bound of $\text{MISE}(\widehat{f}_{t})$. The more standard situation corresponds to the observation of an independent and identically distributed (iid) sample $X_{t,1},\ldots,X_{t,n_{t}}$ but dependent data received also lots of attention (see Rosenblatt, 1971; Nadaraya, 1974; Chanda, 1983; Hart, 1984; Castellana & Leadbetter, 1986; Robinson, 1986; Hall & Hart, 1990; Chesneau, 2014, among others). In the special case of a kernel density estimator
$\widehat{f}_{t}(x)=(n_{t}\,h_{t})^{-1}\sum_{i=1}^{n}K_{d}\left\{h_{t}^{-1}(x-X_{t,i})\right\},$
with $h_{t}$ a bandwidth tending to zero when $n_{t}$ goes to infinity and $K_{d}$ a kernel function, the asymptotic behaviour of $\text{MISE}(\widehat{f}_{t})$ is well known, even when the data $X_{t,1},\ldots,X_{t,n_{t}}$ come from a stationary strong mixing process. When $f_{t}$ is deterministic, Biau (2002) stated that the asymptotic behaviour of $\text{MISE}(\widehat{f}_{t})$ for dependent observations (with arithmetic decay rate for the mixing coefficient) is the same than those obtained for independent data: $\text{MISE}(\widehat{f}_{t})\,=\,O\left(h_{t}^{4}\right)+O\left(n_{t}^{-1}h_{t}^{-1}\right)$. Setting $h_{t}\sim n_{t}^{-1/5}$ in order to balance both terms results in $\text{MISE}(\widehat{f}_{t})\,=\,O\left(n_{t}^{-4/5}\right)$. This result requires that the first two derivatives $f_{t}^{\prime}$ and $f_{t}"$ of the true density function $f_{t}$ are continuous with $\int({f_{t}"})^{2}<\infty$. When $f_{t}$ is a random density, one can write $\text{MISE}(\widehat{f}_{t})\,=\,\text{E}\left\{\text{E}\left(\|\widehat{f}_{t}-f_{t}\|^{2}|f_{t}\right)\right\}$. Then, it is easy to see that the same result holds as soon as
  • (H13)
    $f_{1}^{\prime}$ and $f_{1}"$ are continuous almost surely with $\text{E}\left\{\int\left(f_{1}"\right)^{2}\right\}<\infty$.
The following result simplifies the rate of convergence given in Theorem 1 when the random densities are estimated with the kernel density estimator.
Corollary 1
@@
    $X_{t,1},\ldots,X_{t,n_{t}}$ are iid or generated from a stationary $\rho$-mixing process with arithmetic decay rate for the mixing coefficients,
  • (H17)
    for any $t$, it exists $q>5a/\{4b(1+a)\}$ such that $n_{t}\sim N^{q}$ and $h_{t}\sim N^{-q/5}$,
Then
$\left\|\widehat{m}_{N}(f)\ominus m(f)\right\|_{\mathcal{B}}\,=\,O(h_{r})+O_{P}\left(\left\{N^{a/(1+a)}\,\pi_{f}(h_{r})\right\}^{-1/2}\right).$
(H(H17)) corresponds to a particular case where $\delta_{N}^{2}$, the upper bound of $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)$, is negligible with respect to $N^{-a/(1+a)}$ so that the term $O\left(\delta_{N}\,\pi_{f}(h_{r})^{-1/2}\right)$ vanishes.

### v2

@@ 2.2.3 Probabilistic setting and dependence structure
To better understand the dependence structure of our data, our framework combines two embedded random mechanisms: 1) the random process $\{f\}_{t\in\mathbb{Z}}$ generating the random PDFs $\{f_{1},\ldots,f_{N+1}\}$ and 2) the random process providing for each date $t$ the $n$-tuple $\bm{X}_{t}=\left(X_{t,1},\ldots,X_{t,n}\right)$ for estimating the non-observable random PDF $f_{t}$. The scheme hereafter sums up the overall random situation:
Density-valued process $f_{1}$ $\longrightarrow$ $f_{2}=m(f_{1})\oplus\epsilon_{1}$ $\longrightarrow$ $\cdots$ $\longrightarrow$ $f_{N+1}=m(f_{N})\oplus\epsilon_{N}$
$\big\downarrow$ $\big\downarrow$ $\big\downarrow$
@@
$\big\downarrow$ $\big\downarrow$ $\big\downarrow$
Estimated densities $\widehat{f}_{1}$ $\widehat{f}_{2}$ $\widehat{f}_{N+1}$
Probabilistic setting. Let $(\Omega,\mathcal{A},P)$ be a probability space; the PDF-valued random process $\{f_{t}\}_{t\in\mathbb{Z}}$ is such that, for all $\omega$ in $\Omega$, $f_{t}(\omega,.)$ is in the Bayes Hilbert space $\mathcal{B}_{2}(I)$ and for each $t$, the map $\omega\mapsto f_{t}(\omega,.)$ is a measurable mapping from $(\Omega,\mathcal{A},P)$ to $(\mathcal{B}_{2}(I),\mathcal{G})$ with $\mathcal{G}$ a $\sigma$-algebra. For each $t$, we assume that
  1. 1)
    the PDF of $\bm{X}_{t}$ conditionally to $f_{t}$ is $f_{t}^{n}$, which means that, conditionally to $f_{t}$, $X_{t,1},\ldots,X_{t,n}$ are independently and identically distributed (iid) with PDF of $X_{t,1}$ equals to $f_{t}$,
  2. 2)
    the vector random process $\{\bm{X}_{t}\}_{t\in\mathbb{Z}}$ is independent across $t$ conditionally to the PDF-valued random process $\{f_{t},\}_{t\in\mathbb{Z}}$, which means that, for any finite tuple $(t_{1},\ldots,t_{k})\in\mathbb{Z}^{k}$, the PDF of $(\bm{X}_{t_{1}},\ldots,\bm{X}_{t_{k}})$ conditionally to $\{f_{s}\}_{s\in\mathbb{Z}}$ is equal to $\Pi_{j=1}^{k}\,f_{t_{j}}^{n}$.
Dependence structure. The PDF-valued process $\{f_{t}\}_{t\in\mathbb{Z}}$ is assumed to be a stationary $\rho$-mixing process (see, e.g., Rio, 1993; Bradley, 2005). Set $\mathcal{F}_{p}^{q}=\sigma(f_{t}:\,p\leq t\leq q)$ the $\sigma$-algebra generated by $f_{p},\ldots,f_{q}$ and let $\mathcal{L}^{2}(\mathcal{F}_{p}^{q})$ denote the square-integrable, $\mathcal{F}_{p}^{q}$-measurable rrv space. The quantity
$\rho_{f}(H)=\sup\big\{\left|\text{Corr}(U,\,V)\right|,\,U\in\mathcal{L}^{2}(\mathcal{F}_{-\infty}^{0}),\,V\in\mathcal{L}^{2}(\mathcal{F}_{H}^{+\infty})\big\}$ (4)
is called $\rho$-mixing coefficient of $\{f_{t}\}_{t\in\mathbb{Z}}$; it measures the dependence between $\mathcal{F}_{-\infty}^{0}$ and $\mathcal{F}_{H}^{+\infty}$ in terms of correlation. $\rho_{f}(H)$ is between 0 and 1; the closer this quantity is to 0, the weaker the dependence. By definition, the PDF-valued process $\{f_{t}\}_{t\in\mathbb{Z}}$ is $\rho$-mixing as soon as $\rho_{f}(H)\rightarrow 0$ as $H\rightarrow\infty$. The dependence structure is summarized in the following hypothesis:
  • (H1)
    The random processes $\{f_{t}\}_{t\in\mathbb{Z}}$ and $\{\bm{X}_{t}\}_{t\in\mathbb{Z}}$ are such that:
    • $\bullet$
      $\{f_{t}\}_{t\in\mathbb{Z}}$ is a stationary $\rho$-mixing process,
    • $\bullet$
      for any finite tuple $(t_{1},\ldots,t_{k})\in\mathbb{Z}^{k}$, the PDF of $(\bm{X}_{t_{1}},\ldots,\bm{X}_{t_{k}})$ conditionally to $\{f_{s}\}_{s\in\mathbb{Z}}$ is equal to $\Pi_{j=1}^{k}\,f_{t_{j}}^{n}$.
To ensure the existence of a stationary PDF-valued random process ${f_{t}}_{t\in\mathbb{Z}}$ satisfying Model (1), we require the supplementary hypothesis:
  • (H3)
    It exists $C\in(0,1)$ such that, for all $f$ and $g$ in $\mathcal{B}_{2}(I)$, $\|m(f)\ominus m(g)\|_{\mathcal{B}}\,\leq\,C\,\|f\ominus g\|_{\mathcal{B}}$.
Proposition 1
  • (i)
    Under (H(H3)), there exists a stationary PDF-valued random process that satisfies the nonparametric model (1) and the error process $\{\epsilon_{t}\}_{t\in\mathbb{Z}}$ is necessarily $\rho$-mixing,
  • (ii)
    $\{\bm{X}_{t}\}_{t\in\mathbb{Z}}$ is a stationary $\rho$-mixing process.
The proofs of these results are postponed to the Appendix (see A). There are several reasons to consider (H(H3)). Firstly, Proposition 1 shows the importance of this assumption to prove the existence of a stationary PDF-valued random process consistent with the nonparametric regression model. Secondly, the Lipschitz-type condition is typical in the general nonparametric setting, meaning that the regression is sufficiently smooth. With this assumption, it is possible to control the rate of convergence of the bias (see, e.g., Ferraty et al., 2011, 2012b, for the function-on-function nonparametric regression estimator in the standard space of square integrable functions).
Remark 1
The framework can be extended to accommodate random vectors $\bm{X}_{t}$’s of random lengths $n_{t}$’s, where for each $t$, $\bm{X}_{t}=(X_{t,1},\ldots,X_{t,n_{t}})$ and $n_{t}$ is a random integer. The main asymptotic results given in the next section continue to hold in this more general setting (see Appendix C).
@@ 2.3 Asymptotic properties
This section is devoted to the theoretical properties of our density’s functional kernel estimator. We first introduce the notation and additional assumptions that we need to derive the asymptotic behaviour of $\widehat{m}(f)$.
@@
    There exists a constant $a>1$ such that the $\rho$-mixing coefficient satisfies the condition $\rho_{f}(k)=O(k^{-a})$,
@@
    Let $n=n_{N}$ be a sequence depending on $N$ such that $n$ goes to infinity with $N$; one assumes the existence of a sequence $\delta_{N}$ tending to 0 as $N$ goes to infinity such that, for any $t$, $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|^{2}_{\mathcal{B}}\right)=O(\delta_{N}^{2})$,
@@
    For any deterministic $f$ in $\mathcal{B}_{2}(I)$, the quantity $\pi_{f}(h_{\text{reg}})\coloneqq P\left(\|f_{1}\ominus f\|_{\mathcal{B}}<h_{\text{reg}}\right)$ is strictly positive and it exists $C>0$ and $C^{\prime}>0$, $C\pi_{f}(h_{\text{reg}})\leq\pi_{f}\left(h_{\text{reg}}+o(h_{\text{reg}})\right)\leq C^{\prime}\pi_{f}(h_{\text{reg}})$,
@@
    $h_{\text{reg}}$ tends to 0 with $N$, $N^{a/(1+a)}\,\pi_{f}(h_{\text{reg}})$ goes to infinity with $N$, and it exists $b$ with $1/3<b<1$ such that $\delta_{N}^{2b}=o\left(\pi_{f}(h_{\text{reg}})\right)$ and $\delta_{N}^{1-b}=o\left(h_{\text{reg}}\right)$,
@@
    The kernel function $K_{\text{reg}}$ is Lipschitz and positive on its support [0, 1][0,\,1].
(H(H5)) imposes an arithmetic rate of decay on the mixing coefficients of the PDF-valued random process ${f_{t}}_{t\in\mathbb{Z}}$. This assumption simplifies the convergence rate, although a more general formulation involving the mixing coefficients $\rho_{f}(N)$ can also be derived. (H(H7)) relates the size $n$ of the sample $X_{t,1},\ldots,X_{t,n}$ to the number $N$ of random PDFs. It requires that the mean integrated squared error of the kernel density estimator $\widehat{f}_{t}$, $\text{MISE}(\widehat{f}_{t})\coloneqq\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)$, tends to zero when $n$ (i.e. $N$) goes to infinity. The sequence $\delta_{N}^{2}$ can be interpreted as an upper bound of $\text{MISE}(\widehat{f}_{t})$ in the Bayes Hilbert space. Hypothesis (H(H9)) assumes that the behaviour of the small ball probability $\pi_{f}(h_{\text{reg}})$ is not sensitive to variations in radius negligible with respect to the bandwidth $h_{\text{reg}}$ (see Li & Shao, 2001 for a survey on the small ball probability). According to the isometry property of the clr transformation (see Section 2.1), $\pi_{f}(h_{\text{reg}})=P(\|X-x\|_{2}<h_{\text{reg}})$ with $X=\text{clr}(f_{1})$ and $x=\text{clr}(f)$ in $\mathcal{L}^{2}_{0}(I)$; the class of $\mathcal{L}^{2}_{0}(I)$-valued random functions satisfying (H(H9)) encompasses Gaussian-type random processes (see Ferraty et al., 2019). (H(H11)) relates the asymptotic behaviour of $h_{\text{reg}}$, $\pi_{f}(h_{\text{reg}})$, $N$ and $\delta_{N}^{2}$, the upper bound of $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)$ for any $t$. As $\pi_{f}(h_{\text{reg}})$ tends to zero with $N$ faster than $h_{\text{reg}}$, it is reasonable to require that 2b>1b>02b>1-b>0, which results in $1/3<b<1$. With (H(H13)) and for any $q>0$, $\text{E}\left\{K_{\text{reg}}^{q}(h^{-1}\,\|f_{t}\ominus f\|_{\mathcal{B}})\right\}$ is bounded below and above by a quantity of order $\pi_{f}(h_{\text{reg}})$. For simplicity, we use a basic hypothesis on the kernel function. Nevertheless, it is possible to obtain the more accurate result $\text{E}\left\{K_{\text{reg}}^{q}(h_{\text{reg}}^{-1}\,\|f_{t}\ominus f\|_{\mathcal{B}})\right\}\,=\,C_{f,q}\pi_{f}(h_{\text{reg}})\{1+o(h_{\text{reg}})\}$ as soon as the kernel function $K_{\text{reg}}$ is continuously differentiable on $(0,1)$ with $K_{\text{reg}}(1)>0$, and for all $s\in(0,1)$, $K_{\text{reg}}^{\prime}(s)\leq 0$ and the ratio $\pi_{f}(h_{\text{reg}}\,s)/\pi_{f}(h_{\text{reg}})$ tends to some positive value as $h_{\text{reg}}$ tends to 0.
Once notation and assumptions are introduced, we can provide our main result by giving the pointwise convergence rate of our density-on-density kernel regression estimator.
Theorem 1
If (H(H1))-(H(H13)) are fulfilled, then, for any deterministic $f$ in $\mathcal{B}_{2}(I)$,
$\left\|\widehat{m}_{N}(f)\ominus m(f)\right\|_{\mathcal{B}}\,=\,O(h_{\text{reg}})+O\left(\delta_{N}\,\pi_{f}(h_{\text{reg}})^{-1/2}\right)+O_{P}\left(\left\{N^{a/(1+a)}\,\pi_{f}(h_{\text{reg}})\right\}^{-1/2}\right).$
The first two terms correspond to the bias part. $O(h_{\text{reg}})$, which is not surprising (see the bias part of Theorem 6.11 p. 80 with $\beta=1$ in Ferraty & Vieu, 2006), is derived from the regularity assumption acting on the regression operator $m$ (see (H(H3))). The second term, which is not usual, involves $\delta_{N}$ and the small ball probability $\pi_{f}(h_{\text{reg}})$. This is the price to pay when the random PDFs are not directly observed and are approximated by means of some estimating procedure, where $\delta_{N}$ is the upper bound of the root MISE of $\widehat{f}_{t}$ for any $t$. For the variance part, remember that $a$ controls the asymptotic behaviour of the mixing coefficient of the sequence $\{f_{t}\}_{t\in\mathbb{Z}}$. So, the impact of the dependence model appears directly in the exponent of the sample size $N$ of the random densities. In the situation of a sample of independent explanatory functional random variables, we would get $N$ instead of $N^{a/(1+a)}$ (see the variance part of Theorem 6.11, p. 80 in Ferraty & Vieu, 2006, where the $\log n$ vanishes when considering the convergence in probability).
Remark 2
“The proof of Theorem 1 (see Section B.2 in Appendix B) does not use the fact that the Bayes Hilbert space is isometric to $\mathcal{L}_{0}^{2}(I)$. Consequently, as soon as the unobserved objects can be estimated consistently, Theorem 1 remains valid for non-Euclidean spaces (i.e., spaces not isometric to any Euclidean space).
Let us now focus on the kernel density estimator
$\widehat{f}_{t}(x)=(n\,h_{t})^{-1}\sum_{i=1}^{n}K_{\text{kde}}\left\{h_{t}^{-1}(x-X_{t,i})\right\},$ (5)
with $h_{t}$ a bandwidth tending to zero when $n$ goes to infinity and $K_{\text{kde}}$ is a standard univariate kernel function with support of $K_{\text{kde}}$ equals to [1,1][-1,1] (i.e. symmetric, integrates to 1, and $\int_{-1}^{1}u^{2}\,K_{\text{kde}}(u)\,du<\infty$). When $f_{t}$ is deterministic, the asymptotic approximation of $\text{MISE}(\widehat{f}_{t})$ based on the iid sample $X_{t,1},\ldots,X_{t,n}$ is well known: $\text{MISE}(\widehat{f}_{t})\,=\,O\left(h_{t}^{4}\right)+O\left(n^{-1}h_{t}^{-1}\right)$ (for more details on MISE, see Rosenblatt, 1956; Epanechnikov, 1969; Wegman, 1972; Nadaraya, 1974; Marron & Wand, 1992; Wand & Jones, 1994; Hansen, 2005, among others). This result requires that the second derivative $f_{t}^{\prime\prime}$ of the true density function $f_{t}$ is continuous and satisfies $\int({f_{t}"})^{2}<\infty$. The extension of this result to a random density $f_{t}$ with the probabilistic environment defined in (H(H1)) is obtained thanks to Theorem 2 (see (39) in Appendix B.3); it requires the almost sure (a.s.) version of these regularity assumptions:
@@
    $P\left(\left\{\omega\in\Omega,\,f_{1}^{\prime\prime}(\omega,.)\text{ is continuous and }\int_{I}f_{1}^{\prime\prime}(\omega,.)^{2}<\infty\right\}\right)=1$.
The following result simplifies the rate of convergence given in Theorem 1 when the random densities are estimated with the kernel density estimator (5).
@@
    For any $t$, it exists $q>5a/\{4(1+a)\}$ such that $n\asymp N^{q}$ and $h_{t}\asymp N^{-q/5}$,
  • (i)
    $\text{E}\|\widehat{f}_{t}-f_{t}\|_{\mathcal{B}}^{2}=O\left(N^{-4q/5}\right)$,
  • (ii)
    $\left\|\widehat{m}_{N}(f)\ominus m(f)\right\|_{\mathcal{B}}\,=\,O(h_{\text{reg}})+O_{P}\left(\left\{N^{a/(1+a)}\,\pi_{f}(h_{\text{reg}})\right\}^{-1/2}\right)$,
where $u\asymp v$ means that $u$ and $v$ are equivalent in the sense $u=O(v)$ and $v=O(u)$. (H(H17)) corresponds to a particular case where $\delta_{N}^{2}$, the upper bound of $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)$, is negligible with respect to $N^{-a/(1+a)}$ so that the term $O\left(\delta_{N}\,\pi_{f}(h_{\text{reg}})^{-1/2}\right)$ vanishes. The proof of Theorem 2 is postponed to the Appendix (see Section B.3).